eq. resistance in series

Grade:12th Pass

5 Answers

mohit yadav
10 years ago

R=r1+r2

jhansi adapaka
10 years ago

V/I= R1+R2+R3..........................Rn

Khandavalli Satya Srikanth
10 years ago

Applying Kirchoff,s voltage law

V=V1+V2+V3...Vn

IR=I1RI+I2R2...InRn

Current is same therefore

Req=R1+R2...Rn

VEERU YERRAYYA
10 years ago

Equivalent Resistance in series Resistance Req = R1+R2

RAMAKRISHNA CHARI
10 years ago
Equivalent Resistance in series Resistance is Req=R1+R2+.......................+Rn